OpenStudy (anonymous):

Find the value of the variable(s). If your answer is an integer, leave it in simplest radical form. The triangle is not drawn to scale.

OpenStudy (anonymous):

a) 2 \[b) 12\sqrt{3}\] c) 1/2 \[d) 6\]

OpenStudy (whpalmer4):

\[x^2+6^2=12^2\]by Pythagorean theorem \[x^2 = 12^2-6^2=144-36=108\]\[x=\sqrt{9*12} = \sqrt{9*3*4} = 3*2*\sqrt{3}=6\sqrt{3}\]